# Cutting Through Privacy: A Hyperplane-Based Data Reconstruction Attack in Federated Learning

### v1

@@ Appendix C Modifications for Neural Networks with multiple layers
where $\mathbf{W}^{(3)}\in\mathbb{R}^{C\times m}$, $\mathbf{W}^{(2)}\in\mathbb{R}^{m\times N}$, and $\mathbf{W}^{(1)}\in\mathbb{R}^{N\times d}$. For a fixed neuron $i$, we want to compute the gradient $\frac{\partial{\mathcal{L}_{j}}}{\partial{z^{(1)}_{i}}}$ Using the chain rule, we can express it as:
@@
$\frac{\partial{\mathcal{L}_{j}}}{\partial{b^{(1)}_{i}}}=\sum_{m_{1}}...\sum_{k\in\mathcal{C}}\left(\delta_{k}^{(L)}\prod_{l=1}^{L}w_{m_{l},m_{l-1}}\right)w_{m_{1},i},$ (17)
where $\delta_{k}^{(L)}=\frac{\exp{(z^{(L)}_{k})}}{\sum_{c\in\mathcal{C}}\exp{(z^{(L)}_{c})}}$, and $m_{1},...,m_{L}$ indicate the number of rows in $\mathbf{W}^{(L)}$.
@@
To validate our claim, we conduct experiments on ImageNet using models with 1, 2, 3, and 4 hidden layers. The first layer has $N=1000$, while each subsequent layer consists of 100 neurons. We set the malicious parameters of the input and classification layers as described in App. F, and generate the columns of each additional hidden layer from $\mathcal{N}(0,10^{-6})$. The corresponding bias terms are also scaled by a $10^{-3}$ factor. Table 2 demonstrates that, regardless of the number of hidden layers, the reconstruction accuracy of the attack exhibits only minor fluctuations, confirming that the number of layers does not significantly affect the performance of our attack.
@@ Appendix D Bound on the Number of Rounds
$\left\lceil\log_{\lfloor\frac{N}{n}\rfloor+1}\left(\frac{W}{N\cdot\varepsilon}\right)\right\rceil+1\leq\left\lceil\log_{2}\left(\frac{W}{N\cdot\varepsilon}\right)\right\rceil+1,$
@@
Let us finally give an intuition about how to choose the value for $\varepsilon$. Let $\delta$ be the minimum distance that any two input points have. By standard calculations, one can show that their distance projected to a random direction is at least $\delta/d$ in expectation, where $d$ is the dimensionality of the inputs. Assuming that inputs are normalized to $[0,1]^{d}$, we have that the maximum distance of two inputs is $d$. Hence, using an argument similar to Markov’s inequality, we obtain that the distance after projection of two inputs is at least in the order of $\delta/d^{3}$ with constant probability. Finally, applying a union bound over all pairs of inputs, we can set $\varepsilon=\frac{\delta}{d^{3}n^{2}}$. Hence, the number of rounds is upper bounded in the order of $\log_{2}\left(\frac{Wdn}{N\delta}\right)$.
Appendix E Update Rule for Parallel Search
Algorithm 3 describes how to place the hyperplane in our parallel binary search, proposed in Section 4.4.
Algorithm 3 UpdateHyperplanes
Input: set of search intervals $\mathcal{I}$, model input bias ${\mathbf{b}}^{(1)}$
1:  $M=|\mathcal{I}|$
2:  if $M=1$ then
3:     ${\mathbf{b}}^{(1)}\leftarrow\big{\{}l_{1}+k\frac{u_{1}-l_{1}}{N+1}|k=1,\dots,N\big{\}}$
4:  else
5:     sort $\mathcal{I}$ by descending length $(u_{k}-l_{k})$
7:     for $k=1,...,M$ do
8:        if $k<r$ then
10:           $j\leftarrow j+q+1$
11:        else
13:           $j\leftarrow j+q$
14:        end if
15:     end for
16:  end if
17:  Sort ${\mathbf{b}}^{(1)}$ in ascending order.
18:  Return ${\mathbf{b}}^{(1)}$
Appendix F Attacks Configuration
In our attacks, we used the same initialization strategies for ImageNet and HARUS. Each row of the input weights $\mathbf{W}^{(1)}$ is initialized with identical values drawn from a normal distribution $\mathbf{W}^{(1)}_{i}\sim\mathcal{N}(0,10^{-2})$, while classification weights’ column values are drawn from $\mathbf{W}_{i}^{(2)}\sim\mathcal{N}(0,10^{-2})$. However, the choice of the distribution does not represent a limiting factor, and different distribution can be employed. The classification bias values ${\mathbf{b}}^{(2)}_{i}$ are set to $10^{25}$. We chose a large value for the classification bias according to (10). We did not select a specific value for $\varepsilon$. Instead, we iteratively selected as many bias values as were compatible with the number of communication rounds $T$ available to the attack. Consequently, a total of $NT$ bias values were selected following Alg. 3.
@@ Appendix F Attacks Configuration
To minimize reconstruction errors arising from numerical inaccuracy, the results in the main paper are obtained using double precision. However, additional single precision experiments are presented in App. G.
Appendix G Additional experimental results
G.1 CIFAR-10
@@
We conduct additional experiments on the CIFAR-10 dataset Krizhevsky [2009], evaluating our attack on the same two-layers neural network described in the main paper. For these experiments, our malicious model parameters are initialized according to the setup used in the ImageNet experiments, described in App. F, while for CAH, we set $s=0.95$, according to the setup proposed in [Boenisch et al., 2023]. We perform the attack for batch size in {32, 64, 128, 256, 512, 1024, 2048} and evaluate after {1, 2, 5, 10, 20} rounds. Figure 4a illustrates the percentage of reconstructed images after 20 attack rounds. Our attack achieves perfect reconstruction for 2048 data points, while the baseline recovers less than 5% of the samples. Figure 4b shows the effect of the number communication rounds for $n=2048$. Our attack reaches perfect reconstruction after 20 rounds, while CAH’s performance remains below 5% of accuracy. Finally, Figure 5 provides a comprehensive comparison of the accuracy difference between our attack and the baseline across various batch sizes and communication rounds. CAH initially outperforms our method on image data after a single communication round. However, our attack surpasses the baseline after just two rounds, achieving a remarkable 99% accuracy advantage for $n=2048$ after 20 rounds and confirming the same trends observed on ImageNet and Harus.
@@
G.2 HARUS
@@
We investigate the effect of numerical precision by performing additional HARUS experiments in single precision Table 4 shows the results of both our attack and the baseline. The performance of our attack are comparable with the ones presented in the main paper, showing that our attack’s success does not depend on double precision. However, also in this case numerical instability can prevent us from recovering the full batch, as illustrated by the case of 1024 samples where extending the attack from 20 to 50 rounds did not improve the reconstruction rate.
@@
1 Ours 33.33$\pm$22.17 13.54$\pm$6.31 5.47$\pm$2.34 1.69$\pm$0.81 1.24$\pm$0.96 0.72$\pm$0.91 0.29$\pm$0.27 0.15$\pm$0.06
CAH 51.04$\pm$4.77 23.96$\pm$1.80 10.16$\pm$0.78 4.30$\pm$1.41 1.69$\pm$0.41 0.55$\pm$0.06 0.20$\pm$0.08 0.07$\pm$0.04
2 Ours 97.92$\pm$3.61 95.83$\pm$4.77 86.20$\pm$4.71 71.09$\pm$4.75 48.96$\pm$6.88 25.72$\pm$6.31 9.47$\pm$1.50 2.90$\pm$0.60
CAH 73.96$\pm$9.55 39.06$\pm$3.13 17.19$\pm$3.91 7.29$\pm$2.35 3.45$\pm$0.92 1.11$\pm$0.30 0.42$\pm$0.15 0.12$\pm$0.02
5 Ours 100.0$\pm$0.00 100.0$\pm$0.00 100.0$\pm$0.00 100.0$\pm$0.00 97.01$\pm$0.60 87.73$\pm$2.99 62.50$\pm$3.09 32.10$\pm$2.12
CAH 93.75$\pm$3.13 69.27$\pm$5.02 31.51$\pm$0.45 14.58$\pm$0.81 6.77$\pm$0.11 2.83$\pm$0.35 1.14$\pm$0.12 0.33$\pm$0.06
10 Ours 100.0$\pm$0.00 100.0$\pm$0.00 100.0$\pm$0.00 100.0$\pm$0.00 99.87$\pm$0.23 99.41$\pm$0.34 92.95$\pm$0.79 70.61$\pm$2.26
CAH 96.88$\pm$0.00 82.81$\pm$2.71 45.31$\pm$1.56 22.66$\pm$2.17 10.81$\pm$1.08 4.13$\pm$0.50 1.71$\pm$0.13 0.62$\pm$0.01
20 Ours 100.00$\pm$0.00 100.00$\pm$0.00 100.00$\pm$0.00 100.00$\pm$0.00 99.87$\pm$0.23 99.93$\pm$0.11 99.87$\pm$0.06 94.48$\pm$0.59
CAH 100.00$\pm$0.00 94.27$\pm$0.90 60.94$\pm$2.07 32.81$\pm$1.95 16.15$\pm$2.29 6.87$\pm$1.41 2.99$\pm$0.31 1.04$\pm$0.05
30 Ours 100.00$\pm$0.00 100.00$\pm$0.00 100.00$\pm$0.00 100.00$\pm$0.00 99.87$\pm$0.23 99.93$\pm$0.11 99.93$\pm$0.06 99.04$\pm$0.14
CAH 100.00$\pm$0.00 96.88$\pm$1.56 69.79$\pm$0.45 39.84$\pm$1.41 20.51$\pm$1.19 9.15$\pm$1.48 4.04$\pm$0.10 1.38$\pm$0.08
50 Ours 100.00$\pm$0.00 100.00$\pm$0.00 100.00$\pm$0.00 100.00$\pm$0.00 99.87$\pm$0.23 99.93$\pm$0.11 99.93$\pm$0.05 99.90$\pm$0.13
CAH 100.00$\pm$0.00 97.92$\pm$2.39 79.43$\pm$3.16 50.78$\pm$2.03 28.52$\pm$1.19 13.41$\pm$1.94 6.01$\pm$0.42 2.23$\pm$0.10
G.3 ImageNet
Figure 6 evaluates the difference in reconstruction accuracy between our proposed attack and the baseline across varying batch sizes and communication rounds on ImageNet, confirming our attack’s superior performance over CAH in all multi-round scenarios. The prevalence of red, especially in the upper rows and to the right, clearly demonstrates that our attack generally outperforms the baseline in most settings. This highlights the effectiveness of our approach in leveraging multiple communication rounds and larger batch sizes to improve reconstruction accuracy. The blue cells at the bottom of the heatmap, corresponding to the first communication round, suggest that CAH might have an advantage in single-round attacks, particularly for smaller batch sizes, where the baseline benefits from the initialization. Overall, results confirm the same trends as those on HARUS in the main paper.

### v2

@@ Appendix C Modifications for Neural Networks with multiple layers
where $\mathbf{W}^{(3)}\in\mathbb{R}^{C\times m}$, $\mathbf{W}^{(2)}\in\mathbb{R}^{m\times N}$, and $\mathbf{W}^{(1)}\in\mathbb{R}^{N\times d}$. For a fixed neuron $i$, we want to compute the gradient $\frac{\partial{\mathcal{L}_{j}}}{\partial{b^{(1)}_{i}}}$ . Using the chain rule, we can express it as:
@@
$\frac{\partial{\mathcal{L}_{j}}}{\partial{b^{(1)}_{i}}}=\sum_{m_{1}}...\sum_{k\in\mathcal{C}}\left(\gamma_{k}^{(L)}\prod_{l=1}^{L}w_{m_{l},m_{l-1}}\right)w_{m_{1},i},$ (17)
where $\gamma_{k}^{(L)}=\frac{\exp{(z^{(L)}_{k})}}{\sum_{c\in\mathcal{C}}\exp{(z^{(L)}_{c})}}$, and $m_{1},...,m_{L}$ indicate the number of rows in $\mathbf{W}^{(L)}$.
@@
To validate our claim, we conduct experiments on ImageNet using models with 1, 2, 3, and 4 hidden layers. The first layer has $N=1000$, while each subsequent layer consists of 100 neurons. We set the malicious parameters of the input and classification layers as described in App. E, and generate the columns of each additional hidden layer from $\mathcal{N}(0,10^{-6})$. The corresponding bias terms are also scaled by a $10^{-3}$ factor. Table 2 demonstrates that, regardless of the number of hidden layers, the reconstruction accuracy of the attack exhibits only minor fluctuations, confirming that the number of layers does not significantly affect the performance of our attack.
@@ Appendix D Bound on the Number of Rounds
$\left\lceil\log_{\lfloor\frac{N}{n}\rfloor+1}\left(\frac{W}{N\cdot\varepsilon}\right)\right\rceil+1\leq\left\lceil\log_{2}\left(\frac{W}{N\cdot\varepsilon}\right)\right\rceil+1.$ (18)
@@
Finally, we show how to select the value for $\varepsilon$. Let $\mathcal{B}_{{\mathbf{x}}}=\{{\mathbf{x}}_{1},\dots,{\mathbf{x}}_{n}\}$ be the set of samples that the attacker wants to reconstruct. Assume that inputs have a minimum distance
$\Delta=\min_{\begin{subarray}{c}i,j\\ i\neq j\end{subarray}}\lVert{\mathbf{x}}_{i}-{\mathbf{x}}_{j}\rVert.$ (19)
Draw a random direction vector $\mathbf{v}\sim\mathcal{N}(0,\mathbf{I}_{d\times d})$. The distance of two projections along $\mathbf{v}$ of points in $\mathcal{B}_{{\mathbf{x}}}$ will be the random variable
$\mathrm{p}_{\mathbf{v}}({\mathbf{x}}_{i},{\mathbf{x}}_{j})=\langle{\mathbf{x}}_{i}-{\mathbf{x}}_{j},\mathbf{v}\rangle\sim\mathcal{N}(0,\lVert{\mathbf{x}}_{i}-{\mathbf{x}}_{j}\rVert).$ (20)
Therefore, we can compute:
[…]
$\leq\mathbb{P}\left(q\in\left[\frac{-\varepsilon}{\Delta},\frac{\varepsilon}{\Delta}\right]\right)$ (22)
$\leq\frac{1}{\sqrt{2\pi}}2\frac{\varepsilon}{\Delta}=\sqrt{\frac{2}{\pi}}\frac{\varepsilon}{\Delta},$ (23)
where $q\sim\mathcal{N}\left(0,1\right)$. Here, the last inequality uses the fact that the Gaussian pdf is maximized at 0 and our integration interval is symmetric around 0. Then, applying the union bound, we get:
[…]
$\leq\frac{n^{2}}{2}\sqrt{\frac{2}{\pi}}\frac{\varepsilon}{\Delta}.$ (26)
Hence, by selecting $\varepsilon=\sqrt{2\pi}\frac{\Delta}{n^{2}}\delta$, all the projections will be at least $\varepsilon$-distant with high probability. After substituting the value of $\varepsilon$ in (18), the number of rounds becomes:
$\left\lceil\log_{\lfloor\frac{N}{n}\rfloor+1}\left(\frac{Wn^{2}}{\sqrt{2\pi}N\Delta\delta}\right)\right\rceil+1\leq\left\lceil\log_{2}\left(\frac{Wn^{2}}{\sqrt{2\pi}N\Delta\delta}\right)\right\rceil+1,$ (27)
with probability at least $1-\delta$.
Appendix E Attacks Configuration
In our attacks, we used the same initialization strategies for ImageNet and HARUS. Each row of the input weights $\mathbf{W}^{(1)}$ is initialized with identical values drawn from a normal distribution $\mathbf{W}^{(1)}_{i}\sim\mathcal{N}(0,10^{-2})$, while classification weights’ column values are drawn from $\mathbf{W}_{i}^{(2)}\sim\mathcal{N}(0,10^{-2})$. However, the choice of the distribution does not represent a limiting factor, and different distribution can be employed. The classification bias values ${\mathbf{b}}^{(2)}_{i}$ are set to $10^{25}$. We chose a large value for the classification bias according to (10). We did not select a specific value for $\varepsilon$, instead, we iteratively selected as many bias values as were compatible with the number of communication rounds $T$ available to the attacker. Consequently, a total of $NT$ bias values were selected following Alg. 3.
@@ Appendix E Attacks Configuration
To minimize reconstruction errors arising from numerical inaccuracy, the results in the main paper are obtained using double precision. However, additional single precision experiments are presented in App. F.
Appendix F Additional experimental results
F.1 CIFAR-10
@@
We conduct additional experiments on the CIFAR-10 dataset Krizhevsky [2009], evaluating our attack on the same two-layers neural network described in the main paper. For these experiments, our malicious model parameters are initialized according to the setup used in the ImageNet experiments, described in App. E, while for CAH, we set $s=0.95$, according to the setup proposed in [Boenisch et al., 2023]. We perform the attack for batch size in {32, 64, 128, 256, 512, 1024, 2048} and evaluate after {1, 2, 5, 10, 20} rounds. Figure 4a illustrates the percentage of reconstructed images after 20 attack rounds. Our attack achieves perfect reconstruction for 2048 data points, while the baseline recovers less than 5% of the samples. Figure 4b shows the effect of the number communication rounds for $n=2048$. Our attack reaches perfect reconstruction after 20 rounds, while CAH’s performance remains below 5% of accuracy. Finally, Figure 5 provides a comprehensive comparison of the accuracy difference between our attack and the baseline across various batch sizes and communication rounds. CAH initially outperforms our method on image data after a single communication round. However, our attack surpasses the baseline after just two rounds, achieving a remarkable 99% accuracy advantage for $n=2048$ after 20 rounds and confirming the same trends observed on ImageNet and HARUS.
@@
Multiple Local Steps
In Sec 5.1 we assumed updates based on full-batch gradient computations. Here, we extend our attack to the setting in which clients perform $K>1$ local gradient steps. As discussed in Sec. 7, performing multiple local updates modifies the hyperplanes directions and positions, since $\mathbf{W}^{(1)}$ and ${\mathbf{b}}^{(1)}$ are updated at each step. This complicates the precise isolation of individual data points in the local dataset of the attacked client. To mitigate this effect, we propose a strategy that limits the influence of local updates. In particular, we aim to keep the gradients $\frac{\partial\mathcal{L}}{\partial\mathbf{W}^{(1)}}$ and $\frac{\partial\mathcal{L}}{\partial{\mathbf{b}}^{(1)}}$ small compared to the magnitude of $\mathbf{W}^{(1)}$ and ${\mathbf{b}}^{(1)}$, thereby ensuring that parameter updates remain minimal over multiple steps. We consider a two-layer fully connected neural network, as described in the main paper. From (10) we observe that we can effectively limit the magnitude of the gradients by carefully designing the output weights. Specifically, by assigning a large value $\beta$ to half of the elements in $\mathbf{W}^{(2)}_{i}$ and a slightly smaller value $\beta-\eta$ to the other half (with $\eta$ being a small positive constant), the resulting gradients with respect to the bias $b_{i}^{(1)}$ and the input weights $\mathbf{W}^{(1)}_{i}$ remain small—approximately bounded by $\eta$ and $\eta{\mathbf{x}}_{j}$, respectively. In our implementation, we use a vector $\mathbf{v}\in\mathbb{R}^{C}$, where the first half of its components are set to $10^{5}$, and the second half to $10^{5}-\eta$ with $\eta=10^{-2}$. The vector is then repeated across all the columns in $\mathbf{W}^{(2)}$. To avoid gradient cancellation effects (i.e., $\frac{\partial\mathcal{L}}{\partial{\mathbf{b}}^{(1)}}=0$) in case the training batch contains an equal number of samples from each half of the classes, we add small Gaussian noise $\mathcal{N}(0,10^{-4})$ to each component of $\mathbf{v}$.
In parallel, we also aim to minimize updates to $\mathbf{W}^{(2)}$ itself across local steps. By application of the chain rule, we have $\frac{\partial\mathcal{L}_{j}}{\partial\mathbf{W}^{(2)}}=\frac{\partial\mathcal{L}_{j}}{\partial{\mathbf{z}}^{(2)}}\frac{\partial{\mathbf{z}}^{(2)}}{\partial\mathbf{W}^{(2)}}$, where the first term is given by (12), and the second term is ${\mathbf{z}}^{(1)}=\mathbf{W}^{(1)}{\mathbf{x}}_{j}+{\mathbf{b}}^{(1)}$. As the loss gradient $\frac{\partial\mathcal{L}j}{\partial{\mathbf{z}}^{(2)}}$ depends on classification prediction, and ${\mathbf{b}}^{(1)}$ is modified to perform the binary search (Alg.3), we can only directly control $\mathbf{W}^{(1)}$. Assuming that adjacent pixels in images have similar values, we set weights with alternating signs to drive the dot product $\mathbf{W}^{(1)}_{i}{\mathbf{x}}_{j}$ to be close to zero. Specifically, we sample the even elements of $\mathbf{W}^{(1)}_{i}$ from $U(1,2)$ and for odd ones as $\mathbf{W}^{(1)}_{i}\sim U(-2,-1)$.
Due to the presence of local updates, it is no longer sufficient to determine whether two consecutive hyperplanes isolate a point by directly comparing their corresponding observations for exact equality. Instead, we assess whether the observations are sufficiently similar by projecting the observation associated with the larger bias value onto the other, and computing the Euclidean norm of the resulting projection vector. If this value is smaller than $10^{-4}$, we conclude that the corresponding hyperplanes do not separate any image.
In our experiment, we assume that the target client holds $|S|=1024$ images—here, $|S|$ denotes the dataset size of the client, which differs from the training batch size $n$. We set the number of local steps to $K=|S|/n$, with $n=\{2^{3},2^{4},...,2^{10}\}$, and use a learning rate of $10^{-4}$, a standard choice for training on CIFAR10 to ensure convergence.
The results in Figure 6 demonstrate that increasing the number of local steps makes the attack more challenging; however, our method is still able to recover up to 35% of the images even with local steps. In contrast, the baseline attack (CAH) fails to recover more than 20% of the images under the same setting. It is also worth noting that in our setup we considered more local steps $K$ and larger dataset size than what considered in the baseline Boenisch et al. [2023],as their original setting involves $K=5$ local steps and a smaller dataset size ($|S|=200$).
Figure 6: The percentage of perfectly reconstructed inputs for a two-layer FC NN with $N=1000$ neurons in the first layer. The reconstruction is evaluated after 5050 communication rounds.
Robustness to Noise
To assess the robustness of our attack, we conduct experiments in the presence of noise, which may arise from imperfections in real-world communication channels or from defense mechanisms implemented by the client to defend against privacy attacks, such as local differential privacy[Dwork, 2006].
To improve robustness, the server can design model weights similarly to the strategy proposed to adapt to multiple local steps (check App. F.1), but with a different objective. In that case, the server modifies the classification weights $\mathbf{W}^{(2)}$ to get a small value for $\frac{\partial\mathcal{L}}{\partial\mathbf{W}^{(1)}}$ and $\frac{\partial\mathcal{L}}{\partial{\mathbf{b}}^{(1)}}$ , while here the objective is the opposite: the attacker needs to obtain gradients of large magnitude, so that noise can be filtered out and will not have an impact on the reconstruction quality.
Hence, we define a vector $\mathbf{v}\in\mathbb{R}^{C}$ by setting the first half of its components to $10^{5}$ and the second half to $10^{5}-\eta$, where $\eta=10^{3}$. This vector is then replicated across all columns of $\mathbf{W}^{(2)}$.
Figure 7 illustrates the attack performance across different numbers of rounds and varying Gaussian noise variances. The client holds 1024 images, trains using FedSGD with full-batch updates, and uses a learning rate of $10^{-4}$. As expected, the attack accuracy decreases with increasing noise variance; however, our method remains substantially more effective than the baseline (CAH). While our attack performance degrades gracefully as more noise is added on the client side, CAH is unable to reconstruct any sample for both values of $\sigma$ considered (the two corresponding curves overlap). As a consequence, our attack consistently outperforms the baseline, recovering almost 40% more images.
Figure 7: The percentage of perfectly reconstructed inputs on for two-layer fully connected neural network with $N=1000$ neurons in the first layer, under varying numbers of attack rounds and different noise variances $\sigma$.
F.2 HARUS
@@
We investigate the effect of numerical precision by performing additional HARUS experiments in single precision. Table 4 shows the results of both our attack and the baseline. The performance of our attack are comparable with the ones presented in the main paper, showing that the attack’s success does not depend on double precision. However, also in this case precision errors can prevent us from recovering the full batch, as illustrated by the case of 1024 samples where extending the attack from 20 to 50 rounds does not improve the reconstruction rate.
@@
[…]
F.3 ImageNet
Figure 8 evaluates the difference in reconstruction accuracy between our proposed attack and the baseline across varying batch sizes and communication rounds on ImageNet, confirming our attack’s superior performance over CAH in all multi-round scenarios. The prevalence of red, especially in the upper rows and to the right, clearly demonstrates that our attack generally outperforms the baseline in most settings. This highlights the effectiveness of our approach in leveraging multiple communication rounds and larger batch sizes to improve reconstruction accuracy. The blue cells at the bottom of the heatmap, corresponding to the first communication round, suggest that CAH might have an advantage in single-round attacks, particularly for smaller batch sizes, where the baseline benefits from the multiple-directions initialization. Overall, results confirm the same trends as those on HARUS in the main paper.